# Transition Constrained Bayesian Optimization via Markov Decision Processes

### v2

@@ 4 Solving the planning problem
The planning problem, defined as $\min_{d_{\pi}\in\mathcal{D}}\mathcal{U}(d_{\pi})$, can be thought of as analogous to optimizing an acquisition function in traditional BayesOpt, with the added difficulty of doing it in the space of policies. To solve it, following developments in Hazan et al. [20] and Mutny et al. [39], we use the classical Frank-Wolfe algorithm [26]. It proceeds by decomposing the problem into a series of linear optimization sub-problems. Each linearization results in a policy, and we build a mixture policy consisting of optimal policies for each linearization $\pi_{\operatorname{mix},n}=\{(\alpha_{i},\pi_{i})\}_{i=1}^{n}$, and $\alpha_{i}$ step-sizes of Frank-Wolfe. Conveniently, after the linearization of $\mathcal{U}$ the subproblem on the polytope $\mathcal{D}$ corresponds to an RL problem with reward $\nabla\mathcal{U}$ for which many efficient solvers exist. Namely, for a single mixture component we have,
$d_{\pi_{n+1}}=\operatorname*{arg\,min}_{d\in\mathcal{D}}\sum_{x,a,h}\nabla\mathcal{U}(d_{\pi_{\operatorname{mix},n}})(x,a)d_{h}(x,a)$ (10)
@@ 4.2 Continuous MDPs: Model Predictive Control
With continuous search space, the sub-problem can be solved using continuous RL solvers. However, this can be difficult. The intractable part of the problem is that the distribution $d_{\pi}$ needs to be represented in a computable fashion. We represent the distribution by the sequence of actions taken $\{a_{h}\}_{h=1}^{T}$ with the linear state-space model, $x_{h+1}=Ax_{h}+Ba_{h}$. While this formalism is not as general as it could be, it gives us a tractable sub-problem formulation common to control science scenario [47] that is practical for our experiment and captures a vast array of problems. The optimal set of actions is solved with the following problem, where we state it for the full horizon $H$:
@@
such that $||a_{h}||\leq a_{\max},||x_{h}||\leq 0.5,\text{ and }\ x_{h+1}=Ax_{h}+Ba_{h}$, where the known dynamics serves as constraints. Notice that instead of optimizing over the policy $d_{\pi}$, we directly optimize over the parameterizations of the policy $\{a_{h}\}_{h=1}^{H}$. In fact, this formulation is reminiscent of the model predictive control (MPC) optimization problem. Conceptually, these are the same. The only caveat in our case is that unlike in MPC [17], our objective is non-convex and tends to focus on gathering information rather than stability. Due to the non-convexity in this parameterization, we need to solve it heuristically. We identify a number of useful heuristics to solve this problem in Appendix G.
@@ 5 Experiments
Sections 5.15.3 showcase real-world applications under physical transitions constraints, using the discrete version of the algorithm. Section 5.4 benchmarks against other algorithms in the continuous setting, where we consider the additive transition model of Section 4.2 with $A=B=\mathbf{I}$. We include additional results in Appendix B. For each benchmark, we selected reasonable GP hyper-parameters and fixed them during the optimization. These are summarized in Appendix E.2. As we are interested maximizer identification, in discrete problems, we report the proportion of reruns that succeed at identifying the true maximum. For continuous benchmarks, we report inference regret at each iteration: $\text{Regret}_{t}=f(x_{*})-f(x_{\mu,t})$, where $x_{\mu,t}=\operatorname*{arg\,max}_{x\in\mathcal{X}}\mu_{t}(x)$.
Baselines. We include a naive baseline that greedily optimizes the immediate reward to showcase a method with no planning (Greedy-UCB). Likewise, we create a baseline that replaces the gradient in Eq. (10) with Expected Improvement (MDP-EI) a weak version of planning. In the continuous settings, we compare against truncated SnAKe (TrSnaKe) [15], which minimizes movement distance, and against local search region-constrained BayesOpt or LSR [41] for the same task. We compare two variants for approximating the set of maximizers, one using Thompson Sampling (MDP-BO-TS) and one using Upper Confidence Bound (MDP-BO-UCB).
@@ 5.1 Knorr pyrazole synthesis
The kernel. Schrecker et al. [50] indicates the reaction can be approximately represented by simple kinetics via a differential equation model. We use this information along with techniques for representing linear ODE as constraints in GP fitting [37, 3] to create an approximate ODE kernel $k_{ode}$, of which details are left to Appendix H and may be of independent interest. As the above kernel is only an approximation of the true ODE kernel, which itself is imperfect, we must account for the model mismatch. Therefore, we add a squared exponential term to the kernel to ensure a non-parametric correction, i.e.: $k(\tau_{R},B)=\alpha_{ode}k_{ode}(\tau_{R},B)+\alpha_{rbf}k_{rbf}(\tau_{R},B)$.
@@ 5.2 Monitoring Lake Ypacarai
Samaniego et al. [49] investigated automatic monitoring of Lake Ypacarai, and Folch et al. [13] and Yang et al. [61] benchmarked different BayesOpt algorithms for the task of finding the largest contamination source in the lake. We introduce local transition constraints to this benchmark by creating the lake containing obstacles that limit movement (see Figure 10 in the Appendix). Such obstacles in environmental monitoring may include islands or protected areas for animals. We add initial and final state constraints with the goal to model that the boat has to finish at the maintenance port.
@@ 5.3 Free-electron laser: Transition-driven corruption
Apart from hard constraints, we can apply our framework to state-dependent BayesOpt problems involving transitions. For example, the magnitude of noise $\epsilon$ may depend on the transition. This occurs in systems observing equilibration constraints such as a free-electron laser [31]. Using the simplified simulator of this laser [38], we use our framework to model heteroscedastic noise depending on the difference between the current and next state, $\sigma^{2}(x,x^{\prime})=s(1+w||x-x^{\prime}||_{2})$. By choosing $\mathcal{A}=\mathcal{X}$, we rewrite the problem as $\sigma(s,a)=s(1+w||x-a||_{2})$. The larger the move, the more noisy the observation. This creates a problem, where the BayesOpt needs to balance between informative actions and movement, which can be directly implemented in the objective (8) via the matrix $\mathbf{V}(d_{\pi})=\sum_{x,a\in\mathcal{X}}d_{\pi}(x,a)\frac{1}{\sigma^{2}(x,a)}\Phi(x)\Phi(x)^{\top}+\mathbf{I}$. Figure 3(b) reports the comparison between worst-case stateless BO and our algorithm. Our approach substantially improves performance.
@@ 5.4 Synthetic Benchmarks
We benchmark on a variety of classical BayesOpt problems while imposing local movement constraints and considering both immediate and asynchronous feedback (by introducing an observation delay of 25 iterations). We also include the chemistry SnAr benchmark, from Summit [11], which we treat as asynchronous as per Folch et al. [13]. Results are in Figure 4. In the synchronous setting, we found using the UCB maximizer criteria for MDP-BO yields the best results (c.f. Appendix for details of this variant). We also found that LSR performs very competitively on many benchmarks, frequently matching the performance of MDP-BO. In the asynchronous settings we achieved better results using MDP-BO with Thompson sampling. TrSnAKe baseline appears to be competitive in all synthetic benchmarks as well. However, MDP-BO is more robust having less variance in the chosen paths as seen in the quantiles.
@@ 6 Conclusion
We considered transition-constrained BayesOpt problems arising in physical sciences, such as chemical reactor optimization, that require careful planning to reach any system configuration. Focusing on maximizer identification, we formulated the problem with transition constraints using the framework of Markov decision processes and constructed a tractable algorithm for provably and efficiently solving these problems using dynamic programming or model predictive control sub-routines. This work takes an important step towards the larger application of Bayesian Optimization to real-world problems, as we have shown by strong performance in a large variety of problems with physical transitions. Further work could address the continuous variant of the framework to deal with more general transition dynamics.
@@ Acknowledgments
JPF is funded by EPSRC through the Modern Statistics and Statistical Machine Learning (StatML) CDT (grant no. EP/S023151/1) and by BASF SE, Ludwigshafen am Rhein. RM acknowledges support from the BASF / Royal Academy of Engineering Research Chair in Data-Driven Optimisation. This publication was created as part of NCCR Catalysis (grant number 180544), a National Centre of Competence in Research funded by the Swiss National Science Foundation, and was partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and Innovation Program Grant agreement no. 815943. We would also like to thank Linden Schrecker and Ruby Sedgwick for providing valuable feedback on the project.
@@ Appendix A Visual abstract of the algorithm
Transition Constrained BO Problem FormulationObjective Formulation(a) Need to plan ahead Look-ahead Utilities and averaging out uncertainty (intractable) Hypothesis Testing (b) Variance reduction of a trajectory Unconstrained Linear Bandits (Soare et al. [53]) (frequentist) Adaptive Objective with optimality guarantees for Linear Bandits (Fiez et al. [12]) (frequentist) (c) Acquisition function Acquisition function optimization Direct optimization in the space of trajectories (intractable) (d) Relaxation to space of state-action distributions (e) Solvable by Frank-Wolfe Algorithm Hazan et al. [20], Mutny et al. [39] Convergence to optimal Markovian policy (Mutny et al. [39]) Efficient RL solvers, e.g. Dynamic Programming (f) Iterative Reinforcement Learning sub-problem Adaptive resampling at every iteration (g) Non-Markovian Policies
Figure 5: Visual abstract of the work. In black we show the method presented in this paper, with literature connections shown in blue. In red we show solutions which we did not pursue due to intractability. The problem creates the (a) need to plan ahead. To do this, we take inspiration from hypothesis testing and focus on (b) the variance reduction in a set of maximizers, which leads to our (c) acquisition function. The objective is the same as Fiez et al. [12] introduced in the linear bandits literature from a frequentist perspective. To optimize it, we follow developments in Hazan et al. [20], Mutny et al. [39] by (d) relaxing the acquisition function to the space of state-action distributions and (e) solving the planning problem using the Frank-Wolfe algorithm. This consists of iteratively solving tractable (f) reinforcement learning sub-problems which give us optimal Markov policies. We then apply adaptive resampling to obtain (g) non-Markovian policies.
@@ C.1 Derivation of the Bayesian utility
Suppose there are two elements in $\mathcal{Z}_{\text{simple}}=\{z,x_{f}^{\star}\}$. We will generalize to a composite hypothesis later. In two-element case, the probability of the error in incurred due to select $x^{\prime}$ is:
@@
The randomness here is due to the observations $y=f(X_{\text{new}})+\epsilon$ that are used to fit the estimator $\mu_{T}(x)$. Namely due to $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. Given $f$ (equivalently $\theta$), the distribution of our estimator (namely the posterior mean) is Gaussian. Hence, given $f$:
$\mu_{T}\sim\mathcal{N}((\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}\mathbf{V}_{T}\theta,\sigma^{2}\mathbf{X}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-2}\mathbf{X}),$
@@
This is the posterior over the posterior mean as a function. A posterior over the specific evaluation is $\mu_{T}(z)-\mu_{T}(x_{f}^{\star})\sim\mathcal{N}(\underbrace{\theta^{\top}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}\mathbf{V}_{T}(\Phi(z)-\Phi(x^{\star}_{f}))}_{a},\underbrace{\sigma^{2}(\Phi(z)-\Phi(x^{\star}_{f})^{\top}\mathbf{X}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-2}\mathbf{X}^{\top}(\Phi(z)-\Phi(x^{\star}_{f}))}_{b^{2}})$.
Hence, the probability of exceeding 00, which would constitute a mistake is proportional to:
$P(\mu_{T}(z)-\mu_{T}(x_{f}^{\star})\geq 0|f)\propto\exp\left(-\frac{a_{z}^{2}}{b_{z}^{2}}\right)\frac{1}{b_{z}}\propto\frac{b_{z}^{2}}{a_{z}^{2}}+\log(b_{z}),$
where we applied the fact that $\exp(-x)$ is monotone. Abandoning the $\log$ terms, as they are not dominating the performance among other things because $\log(b_{z})$ is negative for small values of the variance $b_{z}$, we arrive at
$\mathbb{E}_{f\sim\operatorname{GP}}[P(\mu_{T}(z)-\mu_{T}(x_{f}^{\star})\geq 0|f)]\propto\mathbb{E}_{f\sim\operatorname{GP}}\left[\frac{b_{z}^{2}}{a_{z}^{2}}\right],$
which is called the Bayes’ factor and is expected failure rate for the set of potential maximizers $\mathcal{Z}_{\text{simple}}$. The expectation is over the posterior including the evaluations $\mathbf{X}_{t}$ (or prior at the very beginning of the procedure). In fact, we can think of the posterior as being the new prior for the future at any time point.
Now assuming that $\mathcal{Z}$ has more than one additional element, we want to ensure the failure rate is small for all other failure modes, all other hypothesis. Technically this means, we have an alternate hypothesis, which is composite. We take the worst-case perspective as its common with composite hypotheses. In expectation over the prior, we want to minimize:
$\min_{\mathbf{X}_{\text{next}}}\mathbb{E}_{f}\left[\sup_{z\in\mathcal{Z}\setminus\{x_{f}^{\star}\}}P(\mu_{T}(z)-\mu_{T}(x_{f}^{\star})\geq 0|f)\right]\stackrel{{\scriptstyle\cdot}}{{\leq}}\min_{\mathbf{X}_{\text{next}}}\mathbb{E}_{f}\left[\sup_{z\in\mathcal{Z}\setminus\{x_{f}^{\star}\}}\frac{k_{\mathbf{X}_{t}\cup\mathbf{X}_{\text{new}}}(z,x_{f}^{\star})}{(f(z)-f(x_{f}^{\star}))^{2}}\right]$ (14)
@@
$a_{z}^{2}$ $=$ $(\theta^{\top}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}\mathbf{V}_{T}(\Phi(z)-\Phi(x^{\star}_{f}))$
$\geq$ $\theta^{\top}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})(\Phi(z)-\Phi(x^{\star}_{f})))^{2}=(\theta^{\top}(\Phi(z)-\Phi(x^{\star}_{f}))^{2}=(f(z)-f(x^{\star}_{f}))^{2}$
$b_{z}^{2}$ $=$ $\sigma^{2}(\Phi(z)-\Phi(x^{\star}_{f})^{\top}\mathbf{X}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-2}\mathbf{X}^{\top}(\Phi(z)-\Phi(x^{\star}_{f}))$
@@
In the last line we have used the same identity as in Eq. (23).
C.2 Upper-bounding the objective: Eliminating $\mathbb{E}_{f}$
The objective Eq. (14) is intractable due to the expectation of the prior and which involves expectation over the maximum $f(x^{\star}_{f})$, which is known to be very difficult to estimate. Interestingly, the numerator is independent of $f$, and hence the only dependence is through the set $\mathcal{Z}$ as well as the denominator.

### v3

@@ 4 Solving the planning problem
The planning problem, defined as $\min_{d_{\pi}\in\mathcal{D}}\mathcal{U}(d_{\pi})$, can be thought of as analogous to optimizing an acquisition function in traditional BayesOpt, with the added difficulty of doing it in the space of policies. See the bottom half of Figure 5 in Appendix A for a breakdown of the different components of our solution. Following developments in Hazan et al. [34] and Mutný et al. [17], we use the classical Frank-Wolfe algorithm [59]. It proceeds by decomposing the problem into a series of linear optimization sub-problems. Each linearization results in a policy, and we build a mixture policy consisting of optimal policies for each linearization $\pi_{\operatorname{mix},n}=\{(\alpha_{i},\pi_{i})\}_{i=1}^{n}$, and $\alpha_{i}$ step-sizes of Frank-Wolfe. Conveniently, after the linearization of $\mathcal{U}$ the subproblem on the polytope $\mathcal{D}$ corresponds to an RL problem with reward $\nabla\mathcal{U}$ for which many efficient solvers exist. Namely, for a single mixture component we have,
$d_{\pi_{n+1}}=\operatorname*{arg\,min}_{d\in\mathcal{D}}\sum_{x,a,h}\nabla\mathcal{U}(d_{\pi_{\operatorname{mix},n}})(x,a)d_{h}(x,a).$ (10)
@@ 4.2 Continuous MDPs: Model Predictive Control
With continuous search space, the sub-problem can be solved using continuous RL solvers. However, this can be difficult. The intractable part of the problem is that the distribution $d_{\pi}$ needs to be represented in a computable fashion. We represent the distribution by the sequence of actions taken $\{a_{h}\}_{h=1}^{T}$ with the linear state-space model, $x_{h+1}=Ax_{h}+Ba_{h}$. While this formalism is not as general as it could be, it gives us a tractable sub-problem formulation common to control science scenario [61] that is practical for our experiments and captures a vast array of problems. The optimal set of actions is solved with the following problem, where we state it for the full horizon $H$:
@@
such that $||a_{h}||\leq a_{\max},x_{h}\in\mathcal{X},\text{ and }\ x_{h+1}=Ax_{h}+Ba_{h}$, where the known dynamics serves as constraints. Notice that instead of optimizing over the policy $d_{\pi}$, we directly optimize over the parameterizations of the policy $\{a_{h}\}_{h=1}^{H}$. In fact, this formulation is reminiscent of the model predictive control (MPC) optimization problem. Conceptually, these are the same. The only caveat in our case is that unlike in MPC [62], our objective is non-convex and tends to focus on gathering information rather than stability. Due to the non-convexity in this parameterization, we need to solve it heuristically. We identify a number of useful heuristics to solve this problem in Appendix G.
@@ 5 Experiments
Sections 5.15.3 showcase real-world applications under physical transitions constraints, using the discrete version of the algorithm. Section 5.4 benchmarks against other algorithms in the continuous setting, where we consider the additive transition model of Section 4.2 with $A=B=\mathbf{I}$. We include additional results in Appendix B. For each benchmark, we selected reasonable GP hyper-parameters and fixed them during the optimization. These are summarized in Appendix E.2. As we are interested maximizer identification, in discrete problems, we report the proportion of reruns that succeed at identifying the true maximum. For continuous benchmarks, we report inference regret at each iteration: $\text{Regret}_{t}=f(x_{*})-f(x_{\mu,t})$, where $x_{\mu,t}=\operatorname*{arg\,max}_{x\in\mathcal{X}}\mu_{t}(x)$. All statistics reported are over 25 different runs.
Baselines. We include a naive baseline that greedily optimizes the immediate reward to showcase a method with no planning (Greedy-UCB). Likewise, we create a baseline that replaces the gradient in Eq. (10) with Expected Improvement [63] (MDP-EI), a weak version of planning. In the continuous settings, we compare against truncated SnAKe (TrSnaKe) [42], which minimizes movement distance, and against local search region-constrained BayesOpt or LSR [22] for the same task. We compare two variants for approximating the set of maximizers, one using Thompson Sampling (MDP-BO-TS) and one using Upper Confidence Bound (MDP-BO-UCB).
@@ 5.1 Knorr pyrazole synthesis
The kernel. Schrecker et al. [27] indicate the reaction can be approximately represented by simple kinetics via a differential equation model. We use this information along with techniques for representing linear ODE as constraints in GP fitting [65, 66] to create an approximate ODE kernel $k_{ode}$ through the featurization:
$\Phi_{ode}(\tau,B)=(1-\mathcal{S}(B))y^{(1)}(\tau,B)+\mathcal{S}(B)y^{(2)}(\tau,B)$
where $y^{(i)}(\tau,B)$ are equal to:
$\gamma_{i}(B)\left(\frac{\lambda^{(i)}_{2}}{\lambda^{(i)}_{1}-\lambda^{(i)}_{2}}e^{\lambda^{(i)}_{1}\tau}-\frac{\lambda^{(i)}_{1}}{\lambda^{(i)}_{1}-\lambda^{(i)}_{2}}e^{\lambda^{(i)}_{2}\tau}+1\right)$
for $i=1,2$, where $\lambda^{(i)}_{1}$ and $\lambda^{(i)}_{2}$ are eigenvalues of the linearized ODE at different stationary points, $\gamma_{1}(B)=B$, $\gamma_{2}(B)=1-B$, and $\mathcal{S}(x):=(1+e^{-\alpha_{sig}(x-0.5)})^{-1}$ is a sigmoid function. Appendix H holds the details and derivations which may be of independent interest. As the above kernel is only an approximation of the true ODE kernel, which itself is imperfect, we must account for the model mismatch. Therefore, we add a squared exponential term to the kernel to ensure a non-parametric correction, i.e.: $k(\tau,B)=\alpha_{ode}k_{ode}(\tau,B)+\alpha_{rbf}(\tau,B)$.
@@ 5.2 Monitoring Lake Ypacarai
Samaniego et al. [20] investigated automatic monitoring of Lake Ypacarai, and Folch et al. [11] and Yang et al. [40] benchmarked different BayesOpt algorithms for the task of finding the largest contamination source in the lake. We introduce local transition constraints to this benchmark by creating the lake containing obstacles that limit movement (see Figure 12 in the Appendix). Such obstacles in environmental monitoring may include islands or protected areas for animals. We add an initial and final state constraint with the goal of modeling that the boat has to finish at a maintenance port.
@@ 5.3 Free-electron laser: Transition-driven corruption
Apart from hard constraints, we can apply our framework to state-dependent BayesOpt problems involving transitions. For example, the magnitude of noise $\epsilon$ may depend on the transition. This occurs in systems observing equilibration constraints such as a free-electron laser [23]. Using the simplified simulator of this laser [67], we use our framework to model heteroscedastic noise depending on the difference between the current and next state, $\sigma^{2}(x,x^{\prime})=s(1+w||x-x^{\prime}||_{2})$. By choosing $\mathcal{A}=\mathcal{X}$, we rewrite the problem as $\sigma(s,a)=s(1+w||x-a||_{2})$. The larger the move, the more noisy the observation. This creates a problem, where the BayesOpt needs to balance between informative actions and movement, which can be directly implemented in the objective (8) via the matrix $\mathbf{V}(d_{\pi})=\sum_{x,a\in\mathcal{X}}d_{\pi}(x,a)\frac{1}{\sigma^{2}(x,a)}\Phi(x)\Phi(x)^{\top}+\frac{1}{TH}\mathbf{I}$. Figure 3(b) reports the comparison between worst-case stateless BO and our algorithm. Our approach substantially improves performance.
@@ 5.4 Synthetic Benchmarks
We benchmark on a variety of classical BayesOpt problems while imposing local movement constraints and considering both immediate and asynchronous feedback (by introducing an observation delay of 25 iterations). We also include the chemistry SnAr benchmark, from Summit [68], which we treat as asynchronous as per Folch et al. [11]. Results are in Figure 4. In the synchronous setting, we found using the UCB maximizer criteria for MDP-BO yields the best results (c.f. Appendix for details of this variant). We also found that LSR performs very competitively on many benchmarks, frequently matching the performance of MDP-BO. In the asynchronous settings we achieved better results using MDP-BO with Thompson sampling. TrSnAKe baseline appears to be competitive in all synthetic benchmarks as well. However, MDP-BO is more robust having less variance in the chosen paths as seen in the quantiles. It is important to highlight that SnAKe and LSR are specialist heuristic algorithms for local box-constraints, and therefore it is not surprising they perform strongly. Our method can be applied to more general settings and therefore it is very encouraging that MDP-BO is able to match these SOTA algorithms in their specialist domain.
@@ 6 Conclusion
We considered transition-constrained BayesOpt problems arising in physical sciences, such as chemical reactor optimization, that require careful planning to reach any system configuration. Focusing on maximizer identification, we formulated the problem with transition constraints using the framework of Markov decision processes and constructed a tractable algorithm for provably and efficiently solving these problems using dynamic programming or model predictive control sub-routines. We showcased strong empirical performance in a large variety of problems with physical transitions, and achieve state-of-the-art results in classical BayesOpt benchmarks under local movement constraints. This work takes an important step towards the larger application of Bayesian Optimization to real-world problems. Further work could address the continuous variant of the framework to deal with more general transition dynamics, or explore the performance of new objective functions.
@@ Acknowledgments
JPF is funded by EPSRC through the Modern Statistics and Statistical Machine Learning (StatML) CDT (grant no. EP/S023151/1) and by BASF SE, Ludwigshafen am Rhein. RM acknowledges support from the BASF / Royal Academy of Engineering Research Chair in Data-Driven Optimisation. This publication was created as part of NCCR Catalysis (grant number 180544), a National Centre of Competence in Research funded by the Swiss National Science Foundation, and was partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and Innovation Program Grant agreement no. 815943. We would also like to thank Linden Schrecker, Ruby Sedgwick, and Daniel Lengyel for providing valuable feedback on the project.
@@ Appendix A Visual abstract of the algorithm
Transition Constrained BO Problem FormulationObjective Formulation(a) Need to plan ahead Look-ahead Utilities and averaging out uncertainty (intractable) Hypothesis Testing (b) Variance reduction of a trajectory Unconstrained Linear Bandits (Soare et al. [46]) (frequentist) Adaptive Objective with optimality guarantees for Linear Bandits (Fiez et al. [30]) (frequentist) (c) Acquisition function Acquisition function optimization Direct optimization in the space of trajectories (intractable) (d) Relaxation to space of state-action distributions (e) Solvable by Frank-Wolfe Algorithm Hazan et al. [34], Mutný et al. [17] Convergence to optimal Markovian policy (Mutný et al. [17]) Efficient RL solvers, e.g. Dynamic Programming (f) Iterative Reinforcement Learning sub-problem Adaptive resampling at every iteration (g) Non-Markovian Policies
Figure 5: Visual abstract of the work. In black we show the method presented in this paper, with literature connections shown in blue. In red we show solutions which we did not pursue due to intractability. The problem creates the (a) need to plan ahead. To do this, we take inspiration from hypothesis testing and focus on (b) the variance reduction in a set of maximizers, which leads to our (c) acquisition function. The objective is the same as Fiez et al. [30] introduced in the linear bandits literature from a frequentist perspective. To optimize it, we follow developments in Hazan et al. [34], Mutný et al. [17] by (d) relaxing the acquisition function to the space of state-action distributions and (e) solving the planning problem using the Frank-Wolfe algorithm. This consists of iteratively solving tractable (f) reinforcement learning sub-problems which give us optimal Markov policies. We then apply adaptive resampling to obtain (g) non-Markovian policies.
@@ B.4 Computational study
B.5 Median plots for Ypacarai and reactor experiments
In Figures 10 and 11 we give the median and quantile plots for the Knorr pyrazole synthesis and the Ypacarai experiment, which were not included in the main paper to avoid cluttering the graphics.
(a) Episodic feedback results
(b) Immediate feedback results
Figure 10: Median and 10th/90th quantile plots for Knorr pyrazole synthesis experiment.
(a) Episodic feedback results
(b) Immediate feedback results
Figure 11: Median and 10th/90th quantile plots for Ypacarai experiment.
@@ C.1 Derivation of the Bayesian utility
Suppose there are two elements in $\mathcal{Z}_{\text{simple}}=\{z,x_{f}^{\star}\}$. We will generalize to a composite hypothesis later. In two-element case, the probability of the error in incurred due to selecting $z$ is:
@@
. The randomness here is due to the observations $y=f(X_{\text{new}})+\epsilon$ that are used to fit the estimator $\mu_{T}(x)$. Namely due to $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. Given $f$ (equivalently $\theta$), the distribution of our estimator (namely the posterior mean) is Gaussian. Hence, given $f$:
$\mu_{T}\sim\mathcal{N}((\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}\mathbf{V}_{T}\theta,\sigma^{2}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}V_{T}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}),$
@@
This is the posterior over the posterior mean as a function. A posterior over the specific evaluation is $\mu_{T}(z)-\mu_{T}(x_{f}^{\star})\sim\mathcal{N}(\underbrace{\theta^{\top}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}\mathbf{V}_{T}(\Phi(z)-\Phi(x^{\star}_{f}))}_{a},\underbrace{\sigma^{2}(\Phi(z)-\Phi(x^{\star}_{f})^{\top}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}\mathbf{V}_{T}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}(\Phi(z)-\Phi(x^{\star}_{f}))}_{b^{2}})$.
We can now bound the probability of making an error using a Gaussian tail bound inequality:
$\mathbb{P}(\mu_{T}(z)-\mu_{T}(x^{*})\geq 0)=\mathbb{P}(\mu_{T}(z)-\mu_{T}(x^{*})\geq a_{z}+(-a_{z}))\leq e^{-\frac{a^{2}}{2b^{2}}}$
with the caveat that the inequality only holds when the $a_{z}$ is negative. However note that $a_{z}\rightarrow f(z)-f(x^{*})<0$ as $T\rightarrow\infty$ therefore it will hold once $T$ is large enough. From this we can take logarithms and then the expectation across the randomness in the GP:
$\mathbb{E}_{f\sim GP}\left[\log\mathbb{P}(\mu_{T}(z)-\mu_{T}(x^{*})|f)\right]\leq-\frac{1}{2}\mathbb{E}_{f\sim GP}\left[\frac{a^{2}_{z}}{b^{2}_{z}}\right]$
which is called the log Bayes’ factor and is expected log failure rate for the set of potential maximizers $\mathcal{Z}_{\text{simple}}$. The expectation is over the posterior including the evaluations $\mathbf{X}_{t}$ (or prior at the very beginning of the procedure). In fact, we can think of the posterior as being the new prior for the future at any time point. Now assuming that $\mathcal{Z}$ has more than one additional element, we want to ensure the failure rate is small for all other failure modes, all other hypothesis. Technically this means, we have an alternate hypothesis, which is composite – composed of multiple point hypotheses. We take the worst-case perspective as its common with composite hypotheses. In expectation over the prior, we want to minimize:
$\min_{\mathbf{X}_{\text{next}}}\mathbb{E}_{f}\left[\sup_{z\in\mathcal{Z}\setminus\{x_{f}^{\star}\}}\log P(\mu_{T}(z)-\mu_{T}(x_{f}^{\star})\geq 0|f)\right].$ (14)
For moderate to large $T\gg 0$, we can upper bound this objective via elegant argument to yield a very transparent objective:
$\min_{\mathbf{X}_{\text{next}}}\mathbb{E}_{f}\left[\sup_{z\in\mathcal{Z}\setminus\{x_{f}^{\star}\}}\log P(\mu_{T}(z)-\mu_{T}(x_{f}^{\star})\geq 0|f)\right]\stackrel{{\scriptstyle\cdot}}{{\leq}}-\frac{1}{2}\min_{\mathbf{X}_{\text{next}}}\mathbb{E}_{f}\left[\sup_{z\in\mathcal{Z}\setminus\{x_{f}^{\star}\}}\frac{(f(z)-f(x_{f}^{\star}))^{2}}{k_{\mathbf{X}_{t}\cup\mathbf{X}_{\text{new}}}(z,x_{f}^{\star})}\right]$ (15)
@@
$a_{z}^{2}$ $=$ $(\theta^{\top}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}\mathbf{V}_{T}(\Phi(z)-\Phi(x^{\star}_{f})))^{2}$
$=$ $(\theta^{\top}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}}-\mathbf{I}_{\mathcal{H}})(\Phi(z)-\Phi(x^{\star}_{f})))^{2}$
$=$ $(\theta^{\top}(\Phi(z)-\Phi(x^{\star}_{f}))-\theta^{\top}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}(\Phi(z)-\Phi(x^{\star}_{f})))^{2}$
$\stackrel{{\scriptstyle T\gg 0}}{{\approx}}$ $(\theta^{\top}(\Phi(z)-\Phi(x^{\star}_{f})))^{2}=(f(z)-f(x^{\star}_{f}))^{2}$
$b_{z}^{2}$ $=$ $\sigma^{2}(\Phi(z)-\Phi(x^{\star}_{f})^{\top}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}\mathbf{V}_{T}(\mathbf{V}_{T}+\mathbf{I}_{\mathcal{H}})^{-1}(\Phi(z)-\Phi(x^{\star}_{f}))$
@@
In the last line we have used the same identity as in Eq. (25). We will explain how to eliminate the expectation in Section C.2
C.2 Upper-bounding the objective: Eliminating $\mathbb{E}_{f}$ for large $T$.
The objective Eq. (15) is intractable due to the expectation of the prior and which involves expectation over the maximum $f(x^{\star}_{f})$, which is known to be very difficult to estimate. Interestingly, the denominator is independent of $f$ if we adopt the worst-case perspective over the $x^{\star}_{f}$, and hence the only dependence is through the set $\mathcal{Z}$ as well as the denominator. Given all current prior information, we can determine $\mathcal{Z}$, and hence split the expectation. Let us now express